Exercise 2-8 Cost Behavior; High-Low Method [LO2-4, LO2-5] Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 144,000 kilometers during a year, the average operating cost is 12.2 cents per kilometer. If a truck is driven only 96,000 kilometers during a year, the average operating cost increases to 14.8 cents per kilometer. Required: 1. Using the high-low method, estimate the variable and fixed cost elements of the annual cost of the truck operation. (Do not round your intermediate calculations. Round the Variable cost per kilometer to 3 decimal places and Fixed cost answer to nearest whole dollar amount.) Kilometers Total Annual Driven Cost High level of activity Low level of activity Change per kilometer Variable cost per kilometer Fixed cost per year2. Express the variable and fixed costs in the form Y a bX. (Do not round your intermediate calculations. Round the Variable cost per kilometer to 3 decimal places.) 3. If a truck were driven 120,000 kilometers during a year, what total cost would you expect to be incurred? (Do not round intermediate calculations.) Total annual cost
Answer #1
Hoi Chong Transport Ltd.
(A) (B) (A)*(B)
Kilometer Driven Operating Cost per kilometer Total Annual Cost
1) High Level=(A) 144000 0.122 $                    17,568.00
Low Level=(B) 96000 0.148 $                    14,208.00
Difference=(A)-(B) 48000 $                      3,360.00
Formula=(Change in cost/Change in Kilometer driven)
Variable cost per kilometer=($31776/240000)= 0.070 Per Kilometer
Annual Fixed Cost
Total cost at high activity level $              17,568.00
Less: Variable cost=(144000*$.13324) $              10,080.00
Fixed Cost per year=($17568-$10080) $                7,488.00
2) Equation=Y=a+b X
Y=7488+.070 X
3) If Truck are driven 120000 kilometers during the year , total cost=$7488+.070 X
X=120000
Total cost=$7488+(120000*$.070) $              15,888.00